Project C0 C1 C2
A $ 2,950 + $ 2,850 + $ 2,050
B 2,950 + 2,290 + 1,898

a. Calculate the profitability index for A and B assuming a 23% opportunity cost of capital. (Do not round intermediate calculations. Round your answers to 4 decimal places.)

b. According to the profitability index rule, which project(s) should you accept?

  • Project A

  • Project B

  • Both

  • Neither
Answer #1

a)

Project A:

Profitability index = Present value / initial investment

Present value = 2850 / (1 + 0.23)1 + 2050 / (1 + 0.23)2

Present value = 3,672.0867

Profitability index = 3,672.0867 / 2950

Profitability index = 1.2448

Project B:

Profitability index = Present value / initial investment

Present value = 2290 / (1 + 0.23)1 + 1898 / (1 + 0.23)2

Present value = 3,116.332871

Profitability index = 2,659.729448 / 2500

Profitability index = 1.0564

b)

Both

Projects having profitability index of more than 1 should always be accepted.